OpenStudy (anonymous):

Verify the identity. 4 csc 2x = 2 csc^2x tanx

OpenStudy (anonymous):

\[4 \csc 2x = (2 \csc^2x) (tanx)\]

hartnn (hartnn):

csc = 1/ sin

OpenStudy (anonymous):

\[2(1/\sin(x)) = (1/\sin^2(x)) (tanx / 2)\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

I'm really confused.

OpenStudy (anonymous):

So from beginning to end, what should I write?

hartnn (hartnn):

ok, first left side, 4 csc 2x = 4/ sin 2x write sin 2x = 2 sin x cos x what u get ?

OpenStudy (anonymous):

Should I divide both sides by 2?

OpenStudy (anonymous):

4 / 2 sin x cos x

hartnn (hartnn):

lets work with left side only \(\large \frac{4}{sin 2x}=\frac{2}{sin x.cosx}\) yeah, thats correct, now right side.

hartnn (hartnn):

write csc x =1/sin x tan x = sin x/cos x what u get ?

OpenStudy (anonymous):

Where does the tan x go?

hartnn (hartnn):

tan x was in numerator, replace it by sin x/cos x

OpenStudy (anonymous):

2 csc^2x (sinx / cosx)

hartnn (hartnn):

\(\large 2\csc^2 x . \tan x=2 \frac{1}{sin^2 x}\frac{sin x}{cos x}\) got this ?

hartnn (hartnn):

anything getting cancelled ?

OpenStudy (anonymous):

the sinx.

hartnn (hartnn):

what remains ?

OpenStudy (anonymous):

So 2 (1/cosx)

hartnn (hartnn):

\(\huge 2 \frac{1}{\sin x. \sin x}\frac{\sin x}{\cos x}=\frac{2}{\sin x\cos x }\) got that ?

OpenStudy (anonymous):

Yeah.

hartnn (hartnn):

so, left side and right side both equal to same expression, right ? this proves the identity

OpenStudy (anonymous):

Oh! Wow, thanks so much!

hartnn (hartnn):

your welcome ^_^